OpenStudy (anonymous):

an expression is shown below. which value of x makes the expression equivalent to 10 times the square root of 51? 2 times the square root of 51x

Parth (parthkohli):

\[2\sqrt{51x} = 10\sqrt{51}\]Any ideas?

OpenStudy (anonymous):

a) 5 b) 25 c) 50 d)100

Parth (parthkohli):

\[10\sqrt{51} = 2\times 5\sqrt{51}\]And it's given that \(2\sqrt{51x} = 2\times 5\sqrt{51}\). Begin with cancelling \(2\) and \(\sqrt{51}\) from both sides. You are left with \(\sqrt{x} = 5\)

Parth (parthkohli):

Do you get what I said?

OpenStudy (anonymous):

Yes I do. thank you so much(:

Parth (parthkohli):

So what do you get for the answer?

OpenStudy (anonymous):

a...right?

Parth (parthkohli):

How do you get that?

OpenStudy (anonymous):

..uh yeah im lost Cx im not good with algebra. could you explain?

Parth (parthkohli):

Can you first show what you tried at least?

OpenStudy (anonymous):

I mean like, I don't know it. I don't know what to do at all. I just thought 5 was the answer..but I have no idea how to show the work

Parth (parthkohli):

Did you understand what I did there...?

OpenStudy (anonymous):

no

Parth (parthkohli):

You earlier said you did. Erm

Parth (parthkohli):

Keep reading it. You'll certainly have that "ah!" moment if you try to.

OpenStudy (anonymous):

I have no clue what to do. Could you POSSIBLY explain it to me?

Parth (parthkohli):

OK. Do you see how \(10\sqrt{51} = 2\times 5 \sqrt{51}\)?

OpenStudy (anonymous):

Yes. That parts simple

Parth (parthkohli):

And you want to solve the equation \(10\sqrt{51} = 2\sqrt{51x}\) right?

OpenStudy (anonymous):

Yes.

Parth (parthkohli):

So technically you are solving \(2 \times 5\sqrt{51} = 2\sqrt{51x}\). Got it?

OpenStudy (anonymous):

yeah. but why does 51 stay the same?

Parth (parthkohli):

Why wouldn't it?

OpenStudy (anonymous):

I thought that since 51 is multiplied by x that it would change. but clearly it doesn't. i really do not know.

Parth (parthkohli):

Umm, we're gonna proceed with stuff. But do you understand it till \(2\times 5\sqrt{51}=2\sqrt{51x}\)?

OpenStudy (anonymous):

I don't understand why x cancels out. So no.

Parth (parthkohli):

we haven't done any cancellations till now.

OpenStudy (anonymous):

In the first part of the equation you just put, where is the x? I don't understand how it's equal to the second part.

Parth (parthkohli):

Err... let's do that again. The question wants such a value of \(x\) in \(2\sqrt{51x}\) such that it equals \(10\sqrt{51}\). What would you do here?

OpenStudy (anonymous):

Okay. Nevermind. I don't get it.. ill find help somewhere else.

Parth (parthkohli):

If you want help, why not get it here?